OpenStudy (anonymous):

Use a functions to solve: Mary-Jo and Alixx went out shopping. They were planning on buying 6 shirts each. When they arrived at Love Culture, there was a sale, "Buy 2, Get 11/2 Off". The price per shirt is $17. In total, how much will the girls be spending?

OpenStudy (radar):

I take that to be buy one and another for 1/2 off.

OpenStudy (radar):

Each girl is going to purchase 6 shirts. Given 3 are going to cost $17.00 + 3 at $8.50 Cost for each girl is $51.00 + 25.50 for a total $76.50 each Together they spent $153.00 Now to make a function that would get the same results.

OpenStudy (radar):

f(x)=6(x) +6(x/2) =6(x+x/2) f(x) when x=$17